Calculate the density of a penny using a thickness of 1.52 mm, a diameter of 19.05 mm, the mass of a penny, and Equations 2.7 and 2.11.

Equations 2.7: volume= πr^2h

Equations 2.11: density=  mass/volume
Answer #1

Solution:

The volume of a penny having thichness 1.52 mm is calculated by equation,

V = πr^2h

Given,

h = 1.52 mm

radius(r) = diameter/2 = 19.05/2 = 9.525 mm

Thus,

V = 22/7 x (9.525 mm)^2 x 1.52 mm = 433.409 mm^3

V = 433.409 x 0.001 = 0.4334 cm^3

The mass of a penny = 2.50 g

Hence,

Density = mass./ volume

= 2.50 g / 0.4334 cm^3 = 5.77 g/cm^3